A student adds 200.0 mL of distilled water to 20.0 mL of 4.0 M sodium chloride solution.  

 

What is the concentration of the diluted solution?



1
Expert's answer
2021-09-07T02:07:22-0400

Molar mass of sodium chloride = 58.44 g/mol

= 58.44 × 4 = 233.76

= 233.76/20 = 11.69ml

11.69/233.76 = 0.05M

0.05/0.2L

= 0.25mol/L